# Ditch the Denoiser: Emergence of Noise Robustness in Self-Supervised Learning from Data Curriculum

### v1

@@ 4.4 Probing on Clean Test Set Highlights Better Representation Quality
ImageNet-1k (100 epochs) Gaussian-100 SNR: 4.36 dB N2N + DINOv2 75.8 73.1
DINOv2 w/ NCT 75.2 71.5
@@
Gaussian-255 SNR: 0.72 dB N2N + DINOv2 64.1 57.2
DINOv2 w/ NCT 65.6 54.9
@@
ImageNet-100 (1000 epochs) Gaussian-50 SNR: 8.73 dB N2N + DINOv2 78.5 78.2
@@
Table 3: Linear evaluation performance of applying noise curriculum (NC) to other SSL models on ImageNet-100. Performance is improved for all tested models, highlighting the broader applicability of our approach. The accuracy curve over epochs are visualized in Figure 12 in Section B.5.
@@
In previous sections, linear classification accuracies were measured on validation sets that had the same noise distribution as the training set. To further assess representation quality, we evaluate performance on the clean, original ImageNet validation set. Note that N2N + DINOv2 is evaluated without the denoiser because the clean images are already noise-free. Performance on this clean validation set serves as a strong indicator of how well a model captures representations of the "noise-free" real world. Surprisingly, as highlighted by bold numbers in Table 3, DINOv2 w/ NC or NCT outperforms N2N + DINOv2 in most cases, strongly suggesting that our denoiser-free approach learns more robust and generalizable representations than the two-stage denoising pipeline. This can be attributed to the fact that explicit denoising in the two-stage approach inevitably loses some useful information, leading to a degradation in representation quality in the second stage. This finding has important real-world implications: in scenarios where large-scale noisy data is readily available for pretraining, but only a limited amount of clean data is available for specific downstream tasks, DINOv2 w/ NC and NCT offer more effective solutions for learning transferable representations.
@@ 4.5 Instance Recognition Validates Versatility
To further validate the versatility of our method, we extend the evaluation to another downstream task. Here, we examine the instance-level recognition task on noisy images using the embeddings output by the backbone. The approach is non-parametric as the images’ embeddings are ranked based on their cosine similarity with a query image’s embedding. The results are based on the medium difficulty evaluation setup in the Oxford and Paris datasets, where both easy and hard images are treated as positive images. We measure mean average precision (mAP) and report results in Table 4. We see that the results closely align with Section 4.2. DINOv2 w/ NC greatly outperforms DINOv2 in all comparisons. The mAP of DINOv2 w/ NC is on par with N2N + DINOv2, even at extreme noise levels. For example, at Shot $\lambda=1$ and Speckle $\sigma=255$, the mAP of DINOv2 w/ NC surpasses that of N2N + DINOv2 on the Oxford dataset. DINOv2 w/ NCT outperforms all other methods the majority of the time in high noise settings, again validating its effectiveness.
@@
Gaussian $\sigma$=50 N2N + DINOv2 20.89 38.62
@@
Gaussian $\sigma$=100 N2N + DINOv2 21.83 39.04
@@
Gaussian $\sigma$=255 N2N + DINOv2 15.67 33.68
@@
Shot $\lambda$=10 N2N + DINOv2 21.36 40.00
@@
Shot $\lambda$=3 N2N + DINOv2 21.40 40.39
@@
Shot $\lambda$=1 N2N + DINOv2 18.96 37.22
@@
Speckle $\sigma$=102 N2N + DINOv2 22.01 40.47
@@
Speckle $\sigma$=178.5 N2N + DINOv2 21.31 39.31
@@
Speckle $\sigma$=255 N2N + DINOv2 20.11 39.19
@@ 4.6 DINOv2 w/ NC is Applicable to Diverse SSL Models
To test the applicability of DINOv2 w/ NC beyond DINOv2, we extend the DINOv2 w/ NC to various other SSL models on ImageNet-100, including SimCLR [9], MoCo v3 [11], SimSiam [10], iBOT [46], and DINO [7]. These models encompass both contrastive and non-contrastive approaches, as well as ViT-based [13] and CNN-based architectures [17]. All experiments are conducted with Gaussian ($\sigma$=100) noise without denoised regularization. As shown in Table 3, DINOv2 w/ NC improves over the noisy baseline across all models. The improvement is modest or minor for the first three models, i.e., SimCLR, MoCo v3, and SimSiam. The variation in contrastivity and backbone architecture among these models suggests that neither factor is critical to the method’s effectiveness. In contrast, iBOT and DINO (both are precursors to DINOv2) show substantial gains, similar to the improvements observed in DINOv2, where the SSL Model w/ NC closely matches N2N + SSL Model. This indicates that DINOv2 w/ NC has broad applicability, with more pronounced benefits for models sharing architectural and loss-function similarities with DINOv2. We also benchmark the performance of these models on the instance recognition task in Table 8 of Section B.5, which shows exactly identical trends as the linear evaluation, further reaffirming the generality of our approach. Detailed training configurations are provided in the Section A.4.
@@ 5 Limitations, Conclusion and Future Work
In this work, we proposed a fully self-supervised framework for learning noise-robust visual representations. Leveraging an SSL denoiser and the curriculum training strategy enables DINOv2 to improve significantly upon the noisy baseline. While our benchmark focuses on synthetic noise due to the scarcity of large-scale labeled noisy datasets that allows for standard downstream evaluation, this choice enables controlled, reproducible assessment across noise types. One limitation is the assumption that self-supervised denoiser can denoise reasonably well; while it is a fair assumption given the availability of effective SSL denoisers that can address both i.i.d. and structured noises, it remains an open dependency that could affect performance in more challenging settings. Another key direction for future research is to develop automated, adaptive strategies for curriculum schedule, as the current design requires tuning to determine when to shift from denoised to noisy samples during pretraining. More broadly, our methodology can be extended to other modalities like time series data, including noisy audio, EEG signals and financial data; we aim to establish a unified SSL approach for robust representation learning.
@@ Appendix A Technical Details about Experiments
A.1 Details of Toy Experiment in Section 3.2
@@
The results over 17 runs are visualized in Figure 6. We observe that adopting a noise curriculum not only significantly improves accuracy, but also substantially stabilizes performance, as evidenced by the narrower green box compared to the wider orange box. This highlights the dual benefits of curriculum learning in enhancing both robustness and consistency.
@@ A.2 Regularization Strength Details and Ablation
For results in Figure 5, we select optimal regularization strength based on light sweeps visualized in Figure 7. Specifically, $\lambda$ is set to 1.1, 5.0, 2.0 and 4.0 for Gaussian $\sigma=255$, Shot $\lambda=1$, Speckle $\sigma=255$ and Gaussian $\sigma=100$ respectively. Notably, Figure 7 shows that increasing the strength results in a sharp increase in accuracy followed by fluctuatons around a higher plateau, demonstrating the effectiveness of NCT.
In scaling epoch experiment for ImageNet-100 (Table 1(a)), $\lambda$ is set to 1.1, 0.5 and 0.2 for 200, 500 and 1000 epochs respectively. We find longer training duration requires smaller regularization strength. This agrees with our main finding that scaling training can partially mitigate the performance degradation, thus requires less guidance from denoised teacher. In ImageNet-1k experiment (Table 1(b)), $\lambda$ is set to 2.0 and 1.6 for Gaussian $\sigma=100$ and Gaussian $\sigma=255$.
For evaluation on clean set (Table 3), the strengths are set to the same value as their noisy counterparts detailed in the preceding paragraph.
In instance recognition experiments (Table 4), $\lambda=0.2$ for all settings as we find it generally yields better results. For mixed noise-clean data evaluation shown later in Section B.4, $\lambda$ is set to 1.1.
Figure 7: Regularization strength ablation to determine optimal strength for DINOv2 w/ NCT in Figure 5. Most plots show a rapid increase in linear probing accuracy followed by fluctuations around a higher plateau as strength increases. Due to computational constraints, the sweep was not performed with uniform resolution across all settings and may vary across noise types.
@@ A.3 Training Details for Main DINOv2 Experiments
Training Details for Section 4.2. For curriculum-based training, although the model is trained on denoised images for only 140 epochs, the training dynamics in the first stage, such as learning rate scheduling, are configured based on the total training duration of 200 epochs. In the second stage of training on noisy images, the training dynamics are configured for a total duration of 60 epochs to ensure smooth convergence.
Training Details for Section 4.3. In ImageNet-1k experiments (Table 1(b)), the training configuration for DINOv2 and N2N + DINOv2 are set as default, consistent with its original implementation. In the noisy-stage training of DINOv2 w/ NC and DINOv2 w/ NCT, we reduce the linear learning rate warm-up epochs from 10 to 5, and the teacher temperature warm-up epochs from 30 to 15 to accommodate the reduced training duration for noisy adaptation under the constraint of 100 epochs. Intuitively, less warm-up is needed because denoised pre-training provides a good initialization.
Training Details for SimCLR Baseline in Table 1(b). Consistent with DINOv2, we set batch size to 512 and train for 100 epochs during pre-training. We use the open-source PyTorch implementation from AndrewAtanov/simclr-pytorch. Since SimCLR typically requires more epochs during linear evaluation compared to DINOv2, we train the linear head for 90 epochs (112.59k steps) with a batch size of 1024 to ensure fair comparison.
@@
A.4 Training Details for Other SSL Models in Section 4.6
@@
SimCLR. During training, we use a batch size of 100 and LARS optimizer. After square root scaling, the base learning rate is set to $0.750.075\times\sqrt{100}=$ 0.75. During linear evaluation, we use a batch size of 200 and SGD optimizer. After square root scaling, the base learning rate is set to $0.351.6\times\sqrt{200/4096}=$ 0.35.
MoCo v3. During training, we use a batch size of 128 and AdamW optimizer. After square root scaling, the base learning rate is set to $4\times 10^{-4}\times\sqrt{128/1024}=1.414\times 10^{-4}$. During linear evaluation, we use a batch size of 200 and SGD optimizer. After linear scaling, the base learning rate is set to $0.2*(200/256)=0.15625$.
SimSiam. During training, we use a batch size of 100 and SGD optimizer. After square root scaling, the base learning rate is set to $0.093750.15\times\sqrt{100/256}=$ 0.09375. During linear evaluation, we use a batch size of 200 and LARS optimizer. We found the learning rate provided in SimSiam’s codebase is too small for effectively training the linear head. After some experimentation, we set the base learning rate to 2.34375 after the linear scaling $3.0\times(200/256)$.
@@ A.5 Noise Addition Formula
As stated in Section 4.1, we follow ImageNet-C [18] to introduce Gaussian, Shot, and Speckle noises to the images. Let original image be $x$, and normalized image be $\tilde{x}=x/255$.
@@ A.6 Signal-to-noise Ratio Details
All SNR values given are calculated from the raw noisy images without preprocessing, e.g., resizing, crop. We report the average SNR between the training set and the validation set since they have no significant difference. In the instance recognition, we report the average between the Oxford and Paris datasets. Detailed per-set SNR values are provided in Table 6 and Table 6.
@@ B.1 Ablation Unveils Synergy in Curriculum Learning
Figure 8 illustrates that denoised pretraining contributes nearly twice the accuracy improvement over the noisy baseline compared to restarting training. Notably, their effects are not merely additive, as our combined method outperforms the sum of their individual gains, highlighting a synergistic interaction between the two techniques. The synergy arises because, while denoised pretraining provides robust initial weights, the diminished learning rate in later stages limits adaptation to noisy images. Restarting training resets the learning rate, allowing the model to fully adapt to noise and leverage the learned representations effectively.
@@ B.2 Critical Roles of Alignment and Initialization
Unaligned Frozen Teacher. Instead of extracting the weights of the current teacher backbone at the end of the denoised training phase, we retrain a separate backbone on the same denoised dataset to extract teacher weights. This alternative frozen teacher is then used during the noisy training phase. Unlike the original method, this introduces misaligned embeddings between the frozen teacher, the trainable teacher, and the student. At regularization strength 0.2, as shown by the brown line in Figure 9, this setup results in a marginal performance improvement of just 1.1% over the green baseline, significantly lower than the aligned case depicted by the purple line. At regularization strength 1.1, as shown in Figure 10, the strong unaligned teacher instead destabilizes training, dragging accuracy downward as learning proceeds, ultimately underperforming even the DINOv2 w/ NC. These results highlight the seamless integration between the regularization loss and the curriculum training pipeline, where the first stage’s outputs are essential for effectively regularizing the second stage.
Randomly Initialized Backbone. To further investigate the role of denoised pretraining, we replaced the trainable teacher and student backbones with randomly initialized weights during the noisy training phase while keeping the frozen teacher unchanged. At regularization strength 0.2, this configuration yields extremely poor performance, with accuracy dropping below 20%, as shown by the gray line in Figure 9. Notably, at regularization strength 1.1, the training diverges after 30 epochs as the loss explodes, as shown by the discontinuous grey line in  Figure 10. These results highlight the necessity of denoised pretraining, even in the presence of a regularization term, as it provides a crucial foundation for the training process. Without it, the training fails to leverage the frozen teacher’s guidance.

### v2

@@ 4.4 Probing on Clean Test Set Highlights Better Representation Quality
ImageNet-1k
(100 epochs)
Gaussian-100
SNR: 4.36 dB
N2N + DINOv2 75.8 73.1
DINOv2 w/ NCT 75.2 72.1
@@
Gaussian-255
SNR: 0.72 dB
N2N + DINOv2 64.1 57.2
DINOv2 w/ NCT 65.6 55.8
@@
ImageNet-100
(1000 epochs)
Gaussian-50
SNR: 8.73 dB
N2N + DINOv2 78.5 78.2
@@
Table 3: Linear evaluation performance of applying noise curriculum (NC) to other SSL models on ImageNet-100. Performance is improved for all tested models, highlighting the broader applicability of our approach. The accuracy curve over epochs are visualized in Figure˜12 in Section˜B.5.
@@
In previous sections, linear classification accuracies were measured on validation sets that had the same noise distribution as the training set. To further assess representation quality, we evaluate performance on the clean, original ImageNet validation set. Note that N2N + DINOv2 is evaluated without the denoiser because the clean images are already noise-free. Performance on this clean validation set serves as a strong indicator of how well a model captures representations of the "noise-free" real world. Surprisingly, as highlighted by bold numbers in Table˜3, DINOv2 w/ NC or NCT outperforms N2N + DINOv2 in most cases, strongly suggesting that our denoiser-free approach learns more robust and generalizable representations than the two-stage denoising pipeline. This can be attributed to the fact that explicit denoising in the two-stage approach inevitably loses some useful information, leading to a degradation in representation quality in the second stage. This finding has important real-world implications: in scenarios where large-scale noisy data is readily available for pretraining, but only a limited amount of clean data is available for specific downstream tasks, DINOv2 w/ NC and NCT offer more effective solutions for learning transferable representations.
@@ 4.5 Instance Recognition Validates Versatility
To further validate the versatility of our method, we extend the evaluation to another downstream task. Here, we examine the instance-level recognition task on noisy images using the embeddings output by the backbone. The approach is non-parametric as the images’ embeddings are ranked based on their cosine similarity with a query image’s embedding. The results are based on the medium difficulty evaluation setup in the Oxford and Paris datasets, where both easy and hard images are treated as positive images. We measure mean average precision (mAP) and report results in Table˜4. We see that the results closely align with Section˜4.2. DINOv2 w/ NC greatly outperforms DINOv2 in all comparisons. The mAP of DINOv2 w/ NC is on par with N2N + DINOv2, even at extreme noise levels. For example, at Shot $\lambda=1$ and Speckle $\sigma=255$, the mAP of DINOv2 w/ NC surpasses that of N2N + DINOv2 on the Oxford dataset. DINOv2 w/ NCT outperforms all other methods the majority of the time in high noise settings, again validating its effectiveness.
@@
Gaussian
$\sigma$=50
N2N + DINOv2 20.89 38.62
@@
Gaussian
$\sigma$=100
N2N + DINOv2 21.83 39.04
@@
Gaussian
$\sigma$=255
N2N + DINOv2 15.67 33.68
@@
Shot
$\lambda$=10
N2N + DINOv2 21.36 40.00
@@
Shot
$\lambda$=3
N2N + DINOv2 21.40 40.39
@@
Shot
$\lambda$=1
N2N + DINOv2 18.96 37.22
@@
Speckle
$\sigma$=102
N2N + DINOv2 22.01 40.47
@@
Speckle
$\sigma$=178.5
N2N + DINOv2 21.31 39.31
@@
Speckle
$\sigma$=255
N2N + DINOv2 20.11 39.19
@@ 4.6 DINOv2 w/ NC is Applicable to Diverse SSL Models
To test the applicability of DINOv2 w/ NC beyond DINOv2, we extend the DINOv2 w/ NC to various other SSL models on ImageNet-100, including SimCLR [9], MoCo v3 [11], SimSiam [10], iBOT [47], and DINO [7]. These models encompass both contrastive and non-contrastive approaches, as well as ViT-based [13] and CNN-based architectures [18]. All experiments are conducted with Gaussian ($\sigma$=100) noise without denoised regularization. As shown in Table˜3, DINOv2 w/ NC improves over the noisy baseline across all models. The improvement is modest or minor for the first three models, i.e., SimCLR, MoCo v3, and SimSiam. The variation in contrastivity and backbone architecture among these models suggests that neither factor is critical to the method’s effectiveness. In contrast, iBOT and DINO (both are precursors to DINOv2) show substantial gains, similar to the improvements observed in DINOv2, where the SSL Model w/ NC closely matches N2N + SSL Model. This indicates that DINOv2 w/ NC has broad applicability, with more pronounced benefits for models sharing architectural and loss-function similarities with DINOv2. We also benchmark the performance of these models on the instance recognition task in Table˜8 of Section˜B.5, which shows exactly identical trends as the linear evaluation, further reaffirming the generality of our approach. Detailed training configurations are provided in the Section˜A.4.
@@ 5 Limitations, Conclusion and Future Work
In this work, we proposed a fully self-supervised framework for learning noise-robust visual representations. Leveraging an SSL denoiser and the curriculum training strategy enables DINOv2 to improve significantly upon the noisy baseline. While our benchmark focuses on synthetic noise due to the scarcity of large-scale labeled noisy datasets that allows for standard downstream evaluation, this choice enables controlled, reproducible assessment across noise types. One limitation is the assumption that self-supervised denoiser can denoise reasonably well; while it is a fair assumption as we show in Section˜B.6 that even a very weak denoiser (1 training epoch) can still substantially improve accuracy, it remains an open dependency that could affect performance in more challenging settings. Another key direction for future research is to develop automated, adaptive strategies for curriculum schedule, as the current design requires tuning to determine when to shift from denoised to noisy samples during pretraining. More broadly, our methodology can be extended to other modalities like time series data, including noisy audio, EEG signals and financial data; we aim to establish a unified SSL approach for robust representation learning.
@@ Appendix A Technical Details about Experiments
A.1 Details of Toy Experiment in Section˜3.2
@@
The results over 17 runs are visualized in Figure˜6. We observe that adopting a noise curriculum not only significantly improves accuracy, but also substantially stabilizes performance, as evidenced by the narrower green box compared to the wider orange box. This highlights the dual benefits of curriculum learning in enhancing both robustness and consistency.
@@ A.2 Regularization Strength Details and Ablation
For results in Figure˜5, we select optimal regularization strength based on light sweeps visualized in Figure˜7. Specifically, $\lambda$ is set to 1.1, 5.0, 2.0 and 4.0 for Gaussian $\sigma=255$, Shot $\lambda=1$, Speckle $\sigma=255$ and Gaussian $\sigma=100$ respectively. Notably, Figure˜7 shows that increasing the strength results in a sharp increase in accuracy followed by fluctuatons around a higher plateau, demonstrating the effectiveness of NCT.
In scaling epoch experiment for ImageNet-100 (Table˜1(a)), $\lambda$ is set to 1.1, 0.5 and 0.2 for 200, 500 and 1000 epochs respectively. We find longer training duration requires smaller regularization strength. This agrees with our main finding that scaling training can partially mitigate the performance degradation, thus requires less guidance from denoised teacher. In ImageNet-1k experiment (Table˜1(b)), $\lambda$ is set to 2.0 and 1.6 for Gaussian $\sigma=100$ and Gaussian $\sigma=255$.
For evaluation on clean set (Table˜3), the strengths are set to the same value as their noisy counterparts detailed in the preceding paragraph.
In instance recognition experiments (Table˜4), $\lambda=0.2$ for all settings as we find it generally yields better results. For mixed noise-clean data evaluation shown later in Section˜B.4, $\lambda$ is set to 1.1.
Figure 7: Regularization strength ablation to determine optimal strength for DINOv2 w/ NCT in Figure˜5. Most plots show a rapid increase in linear probing accuracy followed by fluctuations around a higher plateau as strength increases. Due to computational constraints, the sweep was not performed with uniform resolution across all settings and may vary across noise types.
@@ A.3 Training Details for Main DINOv2 Experiments
Training Details for Section˜4.2. For curriculum-based training, although the model is trained on denoised images for only 140 epochs, the training dynamics in the first stage, such as learning rate scheduling, are configured based on the total training duration of 200 epochs. In the second stage of training on noisy images, the training dynamics are configured for a total duration of 60 epochs to ensure smooth convergence.
Training Details for Section˜4.3. In ImageNet-1k experiments (Table˜1(b)), the training configuration for DINOv2 and N2N + DINOv2 are set as default, consistent with its original implementation. In the noisy-stage training of DINOv2 w/ NC and DINOv2 w/ NCT, we reduce the linear learning rate warm-up epochs from 10 to 5, and the teacher temperature warm-up epochs from 30 to 15 to accommodate the reduced training duration for noisy adaptation under the constraint of 100 epochs. Intuitively, less warm-up is needed because denoised pre-training provides a good initialization.
Training Details for SimCLR Baseline in Table˜1(b). Consistent with DINOv2, we set batch size to 512 and train for 100 epochs during pre-training. We use the open-source PyTorch implementation from AndrewAtanov/simclr-pytorch. Since SimCLR typically requires more epochs during linear evaluation compared to DINOv2, we train the linear head for 90 epochs (112.59k steps) with a batch size of 1024 to ensure fair comparison.
@@
A.4 Training Details for Other SSL Models in Section˜4.6
@@
SimCLR. During training, we use a batch size of 100 and LARS optimizer. After square root scaling, the base learning rate is set to $0.075\times\sqrt{100}=0.75$. During linear evaluation, we use a batch size of 200 and SGD optimizer. After square root scaling, the base learning rate is set to $1.6\times\sqrt{200/4096}=0.35$.
MoCo v3. During training, we use a batch size of 128 and AdamW optimizer. After square root scaling, the base learning rate is set to $4\times 10^{-4}\times\sqrt{128/1024}=1.414\times 10^{-4}$. During linear evaluation, we use a batch size of 200 and SGD optimizer. After linear scaling, the base learning rate is set to 0.2(200/256)=0.156250.2*(200/256)=0.15625.
SimSiam. During training, we use a batch size of 100 and SGD optimizer. After square root scaling, the base learning rate is set to $0.15\times\sqrt{100/256}=0.09375$. During linear evaluation, we use a batch size of 200 and LARS optimizer. We found the learning rate provided in SimSiam’s codebase is too small for effectively training the linear head. After some experimentation, we set the base learning rate to 2.34375 after the linear scaling $3.0\times(200/256)$.
@@ A.5 Noise Addition Formula
As stated in Section˜4.1, we follow ImageNet-C [19] to introduce Gaussian, Shot, and Speckle noises to the images. Let original image be $x$, and normalized image be $\tilde{x}=x/255$.
@@ A.6 Signal-to-Noise Ratio Details
All SNR values given are calculated from the raw noisy images without preprocessing, e.g., resizing, crop. We report the average SNR between the training set and the validation set since they have no significant difference. In the instance recognition, we report the average between the Oxford and Paris datasets. Detailed per-set SNR values are provided in Table˜6 and Table˜6.
@@ B.1 Ablation Unveils Synergy in Curriculum Learning
Figure˜8 illustrates that denoised pretraining contributes nearly twice the accuracy improvement over the noisy baseline compared to restarting training. Notably, their effects are not merely additive, as our combined method outperforms the sum of their individual gains, highlighting a synergistic interaction between the two techniques. The synergy arises because, while denoised pretraining provides robust initial weights, the diminished learning rate in later stages limits adaptation to noisy images. Restarting training resets the learning rate, allowing the model to fully adapt to noise and leverage the learned representations effectively.
@@ B.2 Critical Roles of Alignment and Initialization
Unaligned Frozen Teacher. Instead of extracting the weights of the current teacher backbone at the end of the denoised training phase, we retrain a separate backbone on the same denoised dataset to extract teacher weights. This alternative frozen teacher is then used during the noisy training phase. Unlike the original method, this introduces misaligned embeddings between the frozen teacher, the trainable teacher, and the student. At regularization strength 0.2, as shown by the brown line in Figure˜9, this setup results in a marginal performance improvement of just 1.1% over the green baseline, significantly lower than the aligned case depicted by the purple line. At regularization strength 1.1, as shown in Figure˜10, the strong unaligned teacher instead destabilizes training, dragging accuracy downward as learning proceeds, ultimately underperforming even the DINOv2 w/ NC. These results highlight the seamless integration between the regularization loss and the curriculum training pipeline, where the first stage’s outputs are essential for effectively regularizing the second stage.
Randomly Initialized Backbone. To further investigate the role of denoised pretraining, we replaced the trainable teacher and student backbones with randomly initialized weights during the noisy training phase while keeping the frozen teacher unchanged. At regularization strength 0.2, this configuration yields extremely poor performance, with accuracy dropping below 20%, as shown by the gray line in Figure˜9. Notably, at regularization strength 1.1, the training diverges after 30 epochs as the loss explodes, as shown by the discontinuous grey line in  Figure˜10. These results highlight the necessity of denoised pretraining, even in the presence of a regularization term, as it provides a crucial foundation for the training process. Without it, the training fails to leverage the frozen teacher’s guidance.